The following data is the input-output tables for different sectors in 

an economy. Find the technology matrix and also the total output against the 

changes in the final demand given:

INDUSTRY X Y FINAL 

DEMAND

X 15 25 40

Y 20 30 50

D=(

50

75)


1
Expert's answer
2021-05-07T10:00:47-0400

The input output coefficient matrix for X and Y is given by-


"A=\begin{bmatrix}
 0.66 &0.04 \\
 0.2 & 0.33
\end{bmatrix}"


Then the Technology matrix( leontiff matrix) is given by-


"I-A=\begin{bmatrix}
 0.34 &-0.04 \\
 -0.2 & 0.67
\end{bmatrix}"


Imagine that the vector of final demands becomes "y = [y_1, y_2] =[50,70]" .Then, to find the corresponding activity levels in "x = [x_1, x_2, x_3]" we must solve the system "(I \u2212 A)x = y." We have-


"\begin{bmatrix}
 0.34 &-0.04 \\
 -0.2 & 0.67
\end{bmatrix}\begin{bmatrix}
 x\\
 y
\end{bmatrix}=\begin{bmatrix}
 50 \\
 70
\end{bmatrix}"



Solving above matrix for x and y and we get-

"x=165.15,y=153.71"